How many ounces of pure gold that costs $850 per ounce must be mixed with 25 ounces of a gold alloy that costs $500 per ounce to make a new alloy that costs $725 per ounce?
Answer by richwmiller(17219)

850*x+500*25=725(25+x)
850*x+12500=18125+725x
850x-725x=18125-12500
125x=5625
x=45 ounces
check
850*45+500*25=725(25+45)
38250+12500=725(70)
50750=50750
ok